\begin{thebibliography}{77}
\expandafter\ifx\csname natexlab\endcsname\relax\def\natexlab#1{#1}\fi

\bibitem[{Aharoni and Goldberg(2017)}]{aharoni-goldberg-2017-morphological}
Roee Aharoni and Yoav Goldberg. 2017.
\newblock Morphological inflection generation with hard monotonic attention.
\newblock In \emph{Proceedings of the 55th Annual Meeting of the Association
  for Computational Linguistics (Volume 1: Long Papers)}, pages 2004--2015,
  Vancouver, Canada. Association for Computational Linguistics.

\bibitem[{Alnajjar et~al.(2019)Alnajjar, Lepp{\"a}nen, Toivonen
  et~al.}]{alnajjar2019no}
Khalid Alnajjar, Leo Lepp{\"a}nen, Hannu Toivonen, et~al. 2019.
\newblock No time like the present: methods for generating colourful and
  factual multilingual news headlines.
\newblock In \emph{Proceedings of the 10th International Conference on
  Computational Creativity}. Association for Computational Creativity.

\bibitem[{Behravan et~al.(2013)Behravan, Hautam{\"a}ki, and
  Kinnunen}]{behravan2013foreign}
Hamid Behravan, Ville Hautam{\"a}ki, and Tomi Kinnunen. 2013.
\newblock Foreign accent detection from spoken finnish using i-vectors.
\newblock In \emph{INTERSPEECH}, volume 2013, page 14th.

\bibitem[{Behravan et~al.(2015)Behravan, Hautam{\"a}ki, and
  Kinnunen}]{behravan2015factors}
Hamid Behravan, Ville Hautam{\"a}ki, and Tomi Kinnunen. 2015.
\newblock Factors affecting i-vector based foreign accent recognition: A case
  study in spoken finnish.
\newblock \emph{Speech Communication}, 66:118--129.

\bibitem[{Bojanowski et~al.(2017)Bojanowski, Grave, Joulin, and
  Mikolov}]{bojanowski2017enriching}
Piotr Bojanowski, Edouard Grave, Armand Joulin, and Tomas Mikolov. 2017.
\newblock Enriching word vectors with subword information.
\newblock \emph{Transactions of the Association for Computational Linguistics},
  5:135--146.

\bibitem[{Creutz(2018)}]{creutz2018open}
Mathias Creutz. 2018.
\newblock Open subtitles paraphrase corpus for six languages.
\newblock In \emph{Proceedings of the Eleventh International Conference on
  Language Resources and Evaluation (LREC 2018)}.

\bibitem[{Creutz and Sj{\"o}blom(2019)}]{creutz2019toward}
Mathias Creutz and Eetu~Eetu Sj{\"o}blom. 2019.
\newblock Toward automatic improvement of language produced by non-native
  language learners.
\newblock In \emph{Proceedings of the 8th Workshop on Natural Language
  Processing for Computer Assisted Language Learning (NLP4CALL 2019), September
  30, Turku Finland}, 164, pages 20--30. Link{\"o}ping University Electronic
  Press.

\bibitem[{Devlin et~al.(2019)Devlin, Chang, Lee, and
  Toutanova}]{devlin-etal-2019-bert}
Jacob Devlin, Ming-Wei Chang, Kenton Lee, and Kristina Toutanova. 2019.
\newblock {BERT}: Pre-training of deep bidirectional transformers for language
  understanding.
\newblock In \emph{Proceedings of the 2019 Conference of the North {A}merican
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies, Volume 1 (Long and Short Papers)}, pages 4171--4186,
  Minneapolis, Minnesota. Association for Computational Linguistics.

\bibitem[{Drobac(2020)}]{1647b325b0f14e76b14c259d9a24c108}
Senka Drobac. 2020.
\newblock \emph{OCR and post-correction of historical newspapers and journals}.
\newblock Ph.D. thesis, University of Helsinki, Finland.

\bibitem[{Drobac and Lind{\'e}n(2020)}]{drobac2020optical}
Senka Drobac and Krister Lind{\'e}n. 2020.
\newblock Optical character recognition with neural networks and
  post-correction with finite state methods.
\newblock \emph{International Journal on Document Analysis and Recognition
  (IJDAR)}, 23(4):279--295.

\bibitem[{Duong et~al.(2020)Duong, H{\"a}m{\"a}l{\"a}inen, and
  Hengchen}]{duong2020unsupervised}
Quan Duong, Mika H{\"a}m{\"a}l{\"a}inen, and Simon Hengchen. 2020.
\newblock An unsupervised method for ocr post-correction and spelling
  normalisation for finnish.
\newblock \emph{arXiv preprint arXiv:2011.03502}.

\bibitem[{Enarvi et~al.(2017)Enarvi, Smit, Virpioja, and
  Kurimo}]{enarvi2017automatic}
Seppo Enarvi, Peter Smit, Sami Virpioja, and Mikko Kurimo. 2017.
\newblock Automatic speech recognition with very large conversational finnish
  and estonian vocabularies.
\newblock \emph{IEEE/ACM Transactions on Audio, Speech, and Language
  Processing}, 25(11):2085--2097.

\bibitem[{Gatt and Krahmer(2018)}]{gatt2018survey}
Albert Gatt and Emiel Krahmer. 2018.
\newblock Survey of the state of the art in natural language generation: Core
  tasks, applications and evaluation.
\newblock \emph{Journal of Artificial Intelligence Research}, 61:65--170.

\bibitem[{Ginter et~al.(2009)Ginter, Suominen, Pyysalo, and
  Salakoski}]{ginter2009combining}
Filip Ginter, Hanna Suominen, Sampo Pyysalo, and Tapio Salakoski. 2009.
\newblock Combining hidden markov models and latent semantic analysis for topic
  segmentation and labeling: Method and clinical application.
\newblock \emph{International journal of medical informatics}, 78(12):e1--e6.

\bibitem[{Gr{\"o}nroos et~al.(2020)Gr{\"o}nroos, Virpioja, and
  Kurimo}]{gronroos2020morfessor}
Stig-Arne Gr{\"o}nroos, Sami Virpioja, and Mikko Kurimo. 2020.
\newblock Morfessor em+ prune: Improved subword segmentation with expectation
  maximization and pruning.
\newblock In \emph{Proceedings of The 12th Language Resources and Evaluation
  Conference}, pages 3944--3953.

\bibitem[{Haapanen and Lepp{\"a}nen(2020)}]{haapanen2020recycling}
Lauri Haapanen and Leo Lepp{\"a}nen. 2020.
\newblock Recycling a genre for news automation: The production of valtteri the
  election bot.
\newblock \emph{AILA Review}, 33(1):67--85.

\bibitem[{H{\"a}m{\"a}l{\"a}inen(2018{\natexlab{a}})}]{hamalainen2018extracting}
Mika H{\"a}m{\"a}l{\"a}inen. 2018{\natexlab{a}}.
\newblock Extracting a semantic database with syntactic relations for finnish
  to boost resources for endangered uralic languages.
\newblock \emph{The Proceedings of Logic and Engineering of Natural Language
  Semantics 15 (LENLS15)}.

\bibitem[{H{\"a}m{\"a}l{\"a}inen(2018{\natexlab{b}})}]{hamalainen2018harnessing}
Mika H{\"a}m{\"a}l{\"a}inen. 2018{\natexlab{b}}.
\newblock Harnessing nlg to create finnish poetry automatically.
\newblock In \emph{Proceedings of the ninth international conference on
  computational creativity}. Association for Computational Creativity (ACC).

\bibitem[{H{\"a}m{\"a}l{\"a}inen(2018{\natexlab{c}})}]{hamalainen2018poem}
Mika H{\"a}m{\"a}l{\"a}inen. 2018{\natexlab{c}}.
\newblock Poem machine-a co-creative nlg web application for poem writing.
\newblock In \emph{The 11th International Conference on Natural Language
  Generation Proceedings of the Conference}. The Association for Computational
  Linguistics.

\bibitem[{H{\"a}m{\"a}l{\"a}inen and
  Alnajjar(2019{\natexlab{a}})}]{hamalainen2019generating}
Mika H{\"a}m{\"a}l{\"a}inen and Khalid Alnajjar. 2019{\natexlab{a}}.
\newblock Generating modern poetry automatically in finnish.
\newblock In \emph{2019 Conference on Empirical Methods in Natural Language
  Processing and 9th International Joint Conference on Natural Language
  Processing Proceedings of the Conference}. The Association for Computational
  Linguistics.

\bibitem[{H{\"a}m{\"a}l{\"a}inen and
  Alnajjar(2019{\natexlab{b}})}]{hamalainen2019let}
Mika H{\"a}m{\"a}l{\"a}inen and Khalid Alnajjar. 2019{\natexlab{b}}.
\newblock Let’s face it. finnish poetry generation with aesthetics and
  framing.
\newblock In \emph{Proceedings of the 12th International Conference on Natural
  Language Generation}, pages 290--300.

\bibitem[{H{\"a}m{\"a}l{\"a}inen and
  Alnajjar(2019{\natexlab{c}})}]{hamalainen2019template}
Mika H{\"a}m{\"a}l{\"a}inen and Khalid Alnajjar. 2019{\natexlab{c}}.
\newblock A template based approach for training nmt for low-resource uralic
  languages-a pilot with finnish.
\newblock In \emph{Proceedings of the 2019 2nd International Conference on
  Algorithms, Computing and Artificial Intelligence}, pages 520--525.

\bibitem[{H{\"a}m{\"a}l{\"a}inen et~al.(2020)H{\"a}m{\"a}l{\"a}inen, Partanen,
  Alnajjar, Rueter, and Poibeau}]{hamalainen2020automatic}
Mika H{\"a}m{\"a}l{\"a}inen, Niko Partanen, Khalid Alnajjar, Jack Rueter, and
  Thierry Poibeau. 2020.
\newblock Automatic dialect adaptation in finnish and its effect on perceived
  creativity.
\newblock In \emph{11th International Conference on Computational Creativity
  (ICCC’20)}. Association for Computational Creativity.

\bibitem[{H{\"a}m{\"a}l{\"a}inen and
  Rueter(2018)}]{hamalainen-rueter-2018-development}
Mika H{\"a}m{\"a}l{\"a}inen and Jack Rueter. 2018.
\newblock Development of an open source natural language generation tool for
  {F}innish.
\newblock In \emph{Proceedings of the Fourth International Workshop on
  Computational Linguistics of Uralic Languages}, pages 51--58, Helsinki,
  Finland. Association for Computational Linguistics.

\bibitem[{Haverinen et~al.(2014)Haverinen, Nyblom, Viljanen, Laippala, Kohonen,
  Missil{\"a}, Ojala, Salakoski, and Ginter}]{haverinen2013tdt}
Katri Haverinen, Jenna Nyblom, Timo Viljanen, Veronika Laippala, Samuel
  Kohonen, Anna Missil{\"a}, Stina Ojala, Tapio Salakoski, and Filip Ginter.
  2014.
\newblock Building the essential resources for {Finnish}: the {Turku Dependency
  Treebank}.
\newblock \emph{Language Resources and Evaluation}, 48:493--531.
\newblock Open access.

\bibitem[{Hengchen et~al.(2018)Hengchen, Kanner, Marjanen, and
  M{\"a}kel{\"a}}]{hengchen2018comparing}
Simon Hengchen, Antti~Olavi Kanner, Jani~Pekka Marjanen, and Eetu
  M{\"a}kel{\"a}. 2018.
\newblock Comparing topic model stability between finnish, swedish, english and
  french.
\newblock In \emph{Digital Humanities in the Nordic Countries}.

\bibitem[{Honnibal et~al.(2020)Honnibal, Montani, Van~Landeghem, and
  Boyd}]{spacy}
Matthew Honnibal, Ines Montani, Sofie Van~Landeghem, and Adriane Boyd. 2020.
\newblock {spaCy: Industrial-strength Natural Language Processing in Python}.
\newblock 10.5281/zenodo.1212303.

\bibitem[{Hurskainen and Tiedemann(2017)}]{hurskainen2017rule}
Arvi Hurskainen and J{\"o}rg Tiedemann. 2017.
\newblock Rule-based machine translation from english to finnish.
\newblock In \emph{Proceedings of the Second Conference on Machine
  Translation}, pages 323--329.

\bibitem[{Hyv{\"o}nen et~al.(2006)Hyv{\"o}nen, Ruotsalo, H{\"a}ggstr{\"o}m,
  Salminen, Junnila, Virkkil{\"a}, Haaramo, M{\"a}kel{\"a}, Kauppinen, and
  Viljanen}]{hyvonen2006culturesampo}
Eero Hyv{\"o}nen, Tuukka Ruotsalo, Thomas H{\"a}ggstr{\"o}m, Mirva Salminen,
  Miikka Junnila, Mikko Virkkil{\"a}, Mikko Haaramo, Eetu M{\"a}kel{\"a}, Tomi
  Kauppinen, and Kim Viljanen. 2006.
\newblock Culturesampo--finnish culture on the semantic web: The vision and
  first results.
\newblock In \emph{Developments in Artificial Intelligence and the Semantic
  Web-Proceedings of the 12th Finnish AI Conference STeP}, pages 26--27.

\bibitem[{Hämäläinen(2019)}]{uralicnlp_2019}
Mika Hämäläinen. 2019.
\newblock {UralicNLP}: An {NLP} library for {U}ralic languages.
\newblock \emph{Journal of Open Source Software}, 4(37):1345.

\bibitem[{Hämäläinen(2021)}]{mika-endangered}
Mika Hämäläinen. 2021.
\newblock Endangered languages are not low-resourced!
\newblock In Mika H{\"a}m{\"a}l{\"a}inen, Niko Partanen, and Khalid Alnajjar,
  editors, \emph{Multilingual Facilitation}. Rootroo Ltd.

\bibitem[{Jauhiainen(2001)}]{jauhiainen-2001-using}
Tommi Jauhiainen. 2001.
\newblock Using existing written language analyzers in understanding natural
  spoken {F}innish.
\newblock In \emph{Proceedings of the 13th Nordic Conference of Computational
  Linguistics ({NODALIDA} 2001)}, Uppsala, Sweden. Department of Linguistics,
  Uppsala University, Sweden.

\bibitem[{Jauhiainen et~al.(2021)Jauhiainen, Jauhiainen, and
  Lindén}]{jauhiainenjauhiainenjauhiainen}
Tommi Jauhiainen, Heidi Jauhiainen, and Krister Lindén. 2021.
\newblock Suomalais-ugrilaiset kielet ja internet -projekti 2013-2019.
\newblock In Mika H{\"a}m{\"a}l{\"a}inen, Niko Partanen, and Khalid Alnajjar,
  editors, \emph{Multilingual Facilitation}. Rootroo Ltd.

\bibitem[{Kanerva et~al.(2018)Kanerva, Ginter, Miekka, Leino, and
  Salakoski}]{udst:turkunlp}
Jenna Kanerva, Filip Ginter, Niko Miekka, Akseli Leino, and Tapio Salakoski.
  2018.
\newblock Turku neural parser pipeline: An end-to-end system for the conll 2018
  shared task.
\newblock In \emph{Proceedings of the CoNLL 2018 Shared Task: Multilingual
  Parsing from Raw Text to Universal Dependencies}. Association for
  Computational Linguistics.

\bibitem[{Kanerva et~al.(2019)Kanerva, R{\"o}nnqvist, Kekki, Salakoski, and
  Ginter}]{kanerva2019template}
Jenna Kanerva, Samuel R{\"o}nnqvist, Riina Kekki, Tapio Salakoski, and Filip
  Ginter. 2019.
\newblock Template-free data-to-text generation of finnish sports news.
\newblock In \emph{Proceedings of the 22nd Nordic Conference on Computational
  Linguistics}, pages 242--252.

\bibitem[{Kettunen(2015)}]{kettunen2015keep}
Kimmo Kettunen. 2015.
\newblock Keep, change or delete? setting up a low resource ocr post-correction
  framework for a digitized old finnish newspaper collection.
\newblock In \emph{Italian Research Conference on Digital Libraries}, pages
  95--103. Springer.

\bibitem[{Koho et~al.(2019)Koho, Ikkala, Leskinen, Tamper, Tuominen, and
  Hyv{\"o}nen}]{koho2019warsampo}
Mikko Koho, Esko Ikkala, Petri Leskinen, Minna Tamper, Jouni Tuominen, and Eero
  Hyv{\"o}nen. 2019.
\newblock Warsampo knowledge graph: Finland in the second world war as linked
  open data.
\newblock \emph{Semantic Web}, (Preprint):1--14.

\bibitem[{Kornai(2013)}]{kornai2013digital}
Andr{\'a}s Kornai. 2013.
\newblock Digital language death.
\newblock \emph{PloS one}, 8(10):e77056.

\bibitem[{Kutuzov et~al.(2017)Kutuzov, Fares, Oepen, and
  Velldal}]{kutuzov2017word}
Andrei Kutuzov, Murhaf Fares, Stephan Oepen, and Erik Velldal. 2017.
\newblock Word vectors, reuse, and replicability: Towards a community
  repository of large-text resources.
\newblock In \emph{Proceedings of the 58th Conference on Simulation and
  Modelling}, pages 271--276. Link{\"o}ping University Electronic Press.

\bibitem[{Laippala and Ginter(2014)}]{laippala2014syntactic}
Veronika Laippala and Filip Ginter. 2014.
\newblock Syntactic n-gram collection from a large-scale corpus of internet
  finnish.
\newblock In \emph{Human Language Technologies-The Baltic Perspective:
  Proceedings of the Sixth International Conference Baltic HLT}, volume 268,
  page 184.

\bibitem[{Leino et~al.(2020)Leino, Leinonen, Singh, Virpioja, and
  Kurimo}]{leino2020finchat}
Katri Leino, Juho Leinonen, Mittul Singh, Sami Virpioja, and Mikko Kurimo.
  2020.
\newblock Finchat: Corpus and evaluation setup for finnish chat conversations
  on everyday topics.
\newblock \emph{Proc. Interspeech 2020}, pages 429--433.

\bibitem[{Lind{\'e}n et~al.(2013)Lind{\'e}n, Axelson, Drobac, Hardwick,
  Kuokkala, Niemi, Pirinen, and Silfverberg}]{linden2013hfst}
Krister Lind{\'e}n, Erik Axelson, Senka Drobac, Sam Hardwick, Juha Kuokkala,
  Jyrki Niemi, Tommi~A Pirinen, and Miikka Silfverberg. 2013.
\newblock {HFST a system for creating NLP tools}.
\newblock In \emph{International Workshop on Systems and Frameworks for
  Computational Morphology}, pages 53--71. Springer.

\bibitem[{Lind{\'e}n and Carlson(2010)}]{linden2010finnwordnet}
Krister Lind{\'e}n and Lauri Carlson. 2010.
\newblock Finnwordnet--finnish wordnet by translation.
\newblock \emph{LexicoNordica--Nordic Journal of Lexicography}, 17:119--140.

\bibitem[{Lind{\'e}n et~al.(2019)Lind{\'e}n, Haltia, Laine, Luukkonen,
  Piitulainen, and V{\"a}is{\"a}nen}]{linden2019finntransframe}
Krister Lind{\'e}n, Heidi Haltia, Antti Laine, Juha Luukkonen, Jussi
  Piitulainen, and Niina V{\"a}is{\"a}nen. 2019.
\newblock Finntransframe: translating frames in the finnframenet project.
\newblock \emph{Language Resources and Evaluation}, 53(1):141--171.

\bibitem[{Lind{\'e}n et~al.(2020)Lind{\'e}n, Jauhiainen, and
  Hardwick}]{linden2020finnsentiment}
Krister Lind{\'e}n, Tommi Jauhiainen, and Sam Hardwick. 2020.
\newblock Finnsentiment--a finnish social media corpus for sentiment polarity
  annotation.
\newblock \emph{arXiv preprint arXiv:2012.02613}.

\bibitem[{Loukasm{\"a}ki and Makkonen(2019)}]{loukasmaki2019eduskunnan}
Petri Loukasm{\"a}ki and Kimmo Makkonen. 2019.
\newblock Eduskunnan t{\"a}ysistunnon puheenaiheet 1999-2014: miten
  k{\"a}sitell{\"a} lda-aihemalleja?
\newblock \emph{Politiikka}, 61(2).

\bibitem[{Luoma et~al.(2020)Luoma, Oinonen, Pyyk{\"o}nen, Laippala, and
  Pyysalo}]{luoma2020broad}
Jouni Luoma, Miika Oinonen, Maria Pyyk{\"o}nen, Veronika Laippala, and Sampo
  Pyysalo. 2020.
\newblock A broad-coverage corpus for finnish named entity recognition.
\newblock In \emph{Proceedings of The 12th Language Resources and Evaluation
  Conference}, pages 4615--4624.

\bibitem[{M{\"a}kel{\"a} et~al.(2020)M{\"a}kel{\"a}, Lagus, Lahti, S{\"a}ily,
  Tolonen, H{\"a}m{\"a}l{\"a}inen, Kaislaniemi, Nevalainen
  et~al.}]{makela2020wrangling}
Eetu M{\"a}kel{\"a}, Krista Lagus, Leo Lahti, Tanja S{\"a}ily, Mikko Tolonen,
  Mika H{\"a}m{\"a}l{\"a}inen, Samuli Kaislaniemi, Terttu Nevalainen, et~al.
  2020.
\newblock Wrangling with non-standard data.
\newblock In \emph{Proceedings of the Digital Humanities in the Nordic
  Countries 5th Conference Riga, Latvia, October 21-23, 2020}. CEUR-WS. org.

\bibitem[{Malmasi and Dras(2014)}]{malmasi2014finnish}
Shervin Malmasi and Mark Dras. 2014.
\newblock Finnish native language identification.
\newblock In \emph{Proceedings of the Australasian Language Technology
  Association Workshop 2014}, pages 139--144.

\bibitem[{Moshagen et~al.(2014)Moshagen, Rueter, Pirinen, Trosterud, and
  Tyers}]{Moshagen2014}
Sjur Moshagen, Jack Rueter, Tommi Pirinen, Trond Trosterud, and Francis~M.
  Tyers. 2014.
\newblock Open-source infrastructures for collaborative work on under-resourced
  languages.
\newblock The LREC 2014 Workshop “CCURL 2014 - Collaboration and Computing
  for Under-Resourced Languages in the Linked Open Data Era”.

\bibitem[{Nicolai and Yarowsky(2019)}]{nicolai-yarowsky-2019-learning}
Garrett Nicolai and David Yarowsky. 2019.
\newblock Learning morphosyntactic analyzers from the {B}ible via iterative
  annotation projection across 26 languages.
\newblock In \emph{Proceedings of the 57th Annual Meeting of the Association
  for Computational Linguistics}, pages 1765--1774, Florence, Italy.
  Association for Computational Linguistics.

\bibitem[{Nyrkk{\"o}(2018)}]{nyrkko2018building}
Seppo Nyrkk{\"o}. 2018.
\newblock Building a finnish som-based ontology concept tagger and harvester.
\newblock In \emph{Proceedings of the Fourth International Workshop on
  Computational Linguistics of Uralic Languages}, pages 18--25.

\bibitem[{{\"O}hman et~al.(2020){\"O}hman, P{\`a}mies, Kajava, and
  Tiedemann}]{ohman2020xed}
Emily {\"O}hman, Marc P{\`a}mies, Kaisla Kajava, and J{\"o}rg Tiedemann. 2020.
\newblock Xed: A multilingual dataset for sentiment analysis and emotion
  detection.
\newblock In \emph{Proceedings of the 28th International Conference on
  Computational Linguistics}, pages 6542--6552.

\bibitem[{Partanen et~al.(2019)Partanen, H{\"a}m{\"a}l{\"a}inen, and
  Alnajjar}]{partanen-etal-2019-dialect}
Niko Partanen, Mika H{\"a}m{\"a}l{\"a}inen, and Khalid Alnajjar. 2019.
\newblock Dialect text normalization to normative standard {F}innish.
\newblock In \emph{Proceedings of the 5th Workshop on Noisy User-generated Text
  (W-NUT 2019)}, pages 141--146, Hong Kong, China. Association for
  Computational Linguistics.

\bibitem[{Pirinen(2019{\natexlab{a}})}]{pirinen-2019-apertium}
Tommi Pirinen. 2019{\natexlab{a}}.
\newblock Apertium-fin-eng{--}rule-based shallow machine translation for {WMT}
  2019 shared task.
\newblock In \emph{Proceedings of the Fourth Conference on Machine Translation
  (Volume 2: Shared Task Papers, Day 1)}, pages 335--341, Florence, Italy.
  Association for Computational Linguistics.

\bibitem[{Pirinen(2015)}]{pirinen2015development}
Tommi~A Pirinen. 2015.
\newblock Development and use of computational morphology of finnish in the
  open source and open science era: Notes on experiences with omorfi
  development.
\newblock \emph{SKY Journal of Linguistics}, 28:381--393.

\bibitem[{Pirinen(2019{\natexlab{b}})}]{pirinen2019neural}
Tommi~A Pirinen. 2019{\natexlab{b}}.
\newblock Neural and rule-based finnish nlp models—expectations, experiments
  and experiences.
\newblock In \emph{Proceedings of the Fifth International Workshop on
  Computational Linguistics for Uralic Languages}, pages 104--114.

\bibitem[{Porjazovski et~al.(2020)Porjazovski, Leinonen, and
  Kurimo}]{porjazovski2020named}
Dejan Porjazovski, Juho Leinonen, and Mikko Kurimo. 2020.
\newblock Named entity recognition for spoken finnish.
\newblock In \emph{Proceedings of the 2nd International Workshop on AI for
  Smart TV Content Production, Access and Delivery}, pages 25--29.

\bibitem[{Qi et~al.(2020)Qi, Zhang, Zhang, Bolton, and
  Manning}]{qi-etal-2020-stanza}
Peng Qi, Yuhao Zhang, Yuhui Zhang, Jason Bolton, and Christopher~D. Manning.
  2020.
\newblock {S}tanza: A python natural language processing toolkit for many human
  languages.
\newblock In \emph{Proceedings of the 58th Annual Meeting of the Association
  for Computational Linguistics: System Demonstrations}, pages 101--108,
  Online. Association for Computational Linguistics.

\bibitem[{Raitio et~al.(2014)Raitio, Lu, Kane, Suni, Vainio, King, and
  Alku}]{raitio2014voice}
Tuomo Raitio, Heng Lu, John Kane, Antti Suni, Martti Vainio, Simon King, and
  Paavo Alku. 2014.
\newblock Voice source modelling using deep neural networks for statistical
  parametric speech synthesis.
\newblock In \emph{2014 22nd European Signal Processing Conference (EUSIPCO)},
  pages 2290--2294. IEEE.

\bibitem[{Raitio et~al.(2008)Raitio, Suni, Vainio, and
  Alku}]{2be1dd96f88b49c5be7f1a06a9f2d99a}
Tuomo Raitio, Antti Suni, Martti Vainio, and Paavo Alku. 2008.
\newblock Hmm-based finnish text-to-speech system utilizing glottal inverse
  filtering.
\newblock In \emph{Interspeech 2008, Brisbane, Australia, September 22-26,
  2008}.

\bibitem[{R{\"a}m{\"o} and Lepp{\"a}nen(2021)}]{ramo2021using}
Miia R{\"a}m{\"o} and Leo Lepp{\"a}nen. 2021.
\newblock Using contextual and cross-lingual word embeddings to improve variety
  in template-based nlg for automated journalism.
\newblock In \emph{Proceedings of the EACL Hackashop on News Media Content
  Analysis and Automated Report Generation}, pages 62--70.

\bibitem[{Reiter(1994)}]{reiter1994has}
Ehud Reiter. 1994.
\newblock Has a consensus nl generation architecture appeared, and is it
  psycholinguistically plausible?
\newblock In \emph{Proceedings of the Seventh International Workshop on Natural
  Language Generation}.

\bibitem[{Ruokolainen et~al.(2019)Ruokolainen, Kauppinen, Silfverberg, and
  Lind{\'e}n}]{ruokolainen2019finnish}
Teemu Ruokolainen, Pekka Kauppinen, Miikka Silfverberg, and Krister Lind{\'e}n.
  2019.
\newblock A finnish news corpus for named entity recognition.
\newblock \emph{Language Resources and Evaluation}, pages 1--26.

\bibitem[{Silfverberg and Hulden(2018)}]{silfverberg-hulden-2018-initial}
Miikka Silfverberg and Mans Hulden. 2018.
\newblock Initial experiments in data-driven morphological analysis for
  {F}innish.
\newblock In \emph{Proceedings of the Fourth International Workshop on
  Computational Linguistics of Uralic Languages}, pages 98--105, Helsinki,
  Finland. Association for Computational Linguistics.

\bibitem[{Silfverberg et~al.(2016)Silfverberg, Ruokolainen, Lind{\'e}n, and
  Kurimo}]{silfverberg2016finnpos}
Miikka Silfverberg, Teemu Ruokolainen, Krister Lind{\'e}n, and Mikko Kurimo.
  2016.
\newblock Finnpos: an open-source morphological tagging and lemmatization
  toolkit for finnish.
\newblock \emph{Language Resources and Evaluation}, 50(4):863--878.

\bibitem[{Silfverberg and Tyers(2019)}]{silfverberg-tyers-2019-data}
Miikka Silfverberg and Francis Tyers. 2019.
\newblock Data-driven morphological analysis for uralic languages.
\newblock In \emph{Proceedings of the Fifth International Workshop on
  Computational Linguistics for Uralic Languages}, pages 1--14, Tartu, Estonia.
  Association for Computational Linguistics.

\bibitem[{Sj{\"o}blom et~al.(2020)Sj{\"o}blom, Creutz, and
  Scherrer}]{sjoblom2020paraphrase}
Eetu Sj{\"o}blom, Mathias Creutz, and Yves Scherrer. 2020.
\newblock Paraphrase generation and evaluation on colloquial-style sentences.
\newblock In \emph{Proceedings of The 12th Language Resources and Evaluation
  Conference}, pages 1814--1822.

\bibitem[{Thomas et~al.(2018)Thomas, Wessman, Ikkala, Tuominen, Koho,
  Hyv{\"o}nen, and Rohiola}]{thomas2018co}
Suzanne~Elizabeth Thomas, Anna Pia~Frederike Wessman, Esko Ikkala, Jouni~Antero
  Tuominen, Mikko Koho, Eero~Antero Hyv{\"o}nen, and Ville Rohiola. 2018.
\newblock (co-) creating a sustainable platform for finland’s archaeological
  chance finds: The story of sualt.
\newblock In \emph{Digital Heritage and Archaeology in Practice}. University
  press of Florida.

\bibitem[{Tiedemann et~al.(2020)Tiedemann, Nieminen, Aulamo, Kanerva, Leino,
  Ginter, and Papula}]{tiedemann-etal-2020-fiskmo}
J{\"o}rg Tiedemann, Tommi Nieminen, Mikko Aulamo, Jenna Kanerva, Akseli Leino,
  Filip Ginter, and Niko Papula. 2020.
\newblock The {FISKM{\"O}} project: Resources and tools for {F}innish-{S}wedish
  machine translation and cross-linguistic research.
\newblock In \emph{Proceedings of the 12th Language Resources and Evaluation
  Conference}, pages 3808--3815, Marseille, France. European Language Resources
  Association.

\bibitem[{Toivanen et~al.(2012)Toivanen, Toivonen, Valitutti, Gross
  et~al.}]{toivanen2012corpus}
Jukka Toivanen, Hannu Toivonen, Alessandro Valitutti, Oskar Gross, et~al. 2012.
\newblock Corpus-based generation of content and form in poetry.
\newblock In \emph{Proceedings of the third international conference on
  computational creativity}. University College Dublin.

\bibitem[{Ul{\v{c}}ar and Robnik-{\v{S}}ikonja(2020)}]{ulvcar2020high}
Matej Ul{\v{c}}ar and Marko Robnik-{\v{S}}ikonja. 2020.
\newblock High quality elmo embeddings for seven less-resourced languages.
\newblock In \emph{Proceedings of The 12th Language Resources and Evaluation
  Conference}, pages 4731--4738.

\bibitem[{Vankka et~al.(2019)Vankka, Myllykoski, Peltonen, and
  Riippa}]{8931724}
Jouko Vankka, Heikki Myllykoski, Tuomas Peltonen, and Ken Riippa. 2019.
\newblock Sentiment analysis of finnish customer reviews.
\newblock In \emph{2019 Sixth International Conference on Social Networks
  Analysis, Management and Security (SNAMS)}, pages 344--350.

\bibitem[{Varjokallio et~al.(2021)Varjokallio, Virpioja, and
  Kurimo}]{varjokallio2021morphologically}
Matti Varjokallio, Sami Virpioja, and Mikko Kurimo. 2021.
\newblock Morphologically motivated word classes for very large vocabulary
  speech recognition of finnish and estonian.
\newblock \emph{Computer Speech \& Language}, 66:101141.

\bibitem[{Virkkunen et~al.(2018)Virkkunen, Simko, Kallio, Vainio
  et~al.}]{virkkunen2018prosodic}
P{\"a}ivi~Johanna Virkkunen, Juraj Simko, Heini~Henriikka Kallio, Martti~Tapani
  Vainio, et~al. 2018.
\newblock Prosodic features of finnish compound words.
\newblock In \emph{Proceedings of the 9th International Conference on Speech
  Prosody 2018}. International Speech Communications Association.

\bibitem[{Virtanen et~al.(2019)Virtanen, Kanerva, Ilo, Luoma, Luotolahti,
  Salakoski, Ginter, and Pyysalo}]{virtanen2019multilingual}
Antti Virtanen, Jenna Kanerva, Rami Ilo, Jouni Luoma, Juhani Luotolahti, Tapio
  Salakoski, Filip Ginter, and Sampo Pyysalo. 2019.
\newblock Multilingual is not enough: Bert for finnish.
\newblock \emph{arXiv preprint arXiv:1912.07076}.

\bibitem[{Šimko et~al.(2020)Šimko, Vainio, and Suni}]{simko2020}
Juraj Šimko, Martti Vainio, and Antti Suni. 2020.
\newblock {Analysis of speech prosody using WaveNet embeddings: The Lombard
  effect}.
\newblock In \emph{Proc. 10th International Conference on Speech Prosody 2020},
  pages 910--914.

\end{thebibliography}

\documentclass[11pt,a4paper]{article}
\usepackage[hyperref]{acl2018}
\usepackage{url}
\aclfinalcopy % Uncomment this line for the final submission
\title{The Current State of Finnish NLP}

\author{Mika Hämäläinen \\
  Faculty of Arts \\
  University of Helsinki \\
  and Rootroo Ltd \\
  {\tt mika.hamalainen@helsinki.fi} \\\And
  Khalid Alnajjar \\
  Faculty of Arts \\
  University of Helsinki \\
  and Rootroo Ltd \\
  {\tt khalid.alnajjar@helsinki.fi} \\}

\date{}

\begin{document}
\maketitle
\begin{abstract}
  There are a lot of tools and resources available for processing Finnish. In this paper, we survey recent papers focusing on Finnish NLP related to many different subcategories of NLP such as parsing, generation, semantics and speech. NLP research is conducted in many different research groups in Finland, and it is frequently the case that NLP tools and models resulting from academic research are made available for others to use on platforms such as Github.
  \begin{center} \textbf{Tiivistelmä} \end{center}
  Suomen kielen koneelliseen käsittelyyn on tarjolla paljon valmiita työkaluja ja resursseja. Tässä artikkelissa tarkastelemme viimeaikoina julkaistuja tieteellisiä artikkeleita, joissa keskitytään suomen kielen kieliteknologiaan. Tarkastelemme kieliteknologian eri alaluokkia, kuten jäsentämistä, tuottamista, semantiikkaa ja puheetta. kieliteknologista tutkimusta tehdään Suomessa monissa eri tutkimusryhmissä, ja usein akateemisen tutkimuksen tuloksena tuotetut kieliteknologian työkalut ja mallit julkaistaan muiden käytettäväksi esimerkiksi Githubissa.
\end{abstract}

\section{Introduction}

There is no doubt that, within the Uralic language family, Finnish is one of the most well-resourced languages in terms of natural language processing (NLP). This has, however, not always been the case. Currently, NLP research conducted for Finnish has started to fragment into research outputs of several different research groups, and there is no survey paper out there that would describe the current state of Finnish NLP.

We hope that this survey paper clarifies the current situation and makes it clearer for people working in the academia outside of Finnish universities or in the industry and also for students. As it has been discussed before \cite{mika-endangered}, Finnish is certainly not a low-resourced language, and our current survey further proves this point.

It is also important for researchers working on other smaller Uralic languages to see what has been done for Finnish in terms of NLP to see what the possible and meaningful directions are for further developing the resources needed. Especially since Uralic language share the same feature of rich morphology, which is something that commonly causes problems for computers.

\section{Finnish NLP}

In this section, we present a survey on the current state of Finnish NLP. We have tried to gather most of the current research on the topic, but we are certain that there are some research out there we have not been able to find. We have categorized the surveyed research outputs into parsing, generation, semantics and speech.

\subsection{Parsing}

Starting from morphology, stemming and spell checking Finnish is well supported in multiple commercial applications such as Microsoft and Google products. In the open-source world, low-level tasks such as stemming and spell checking can be conducted with Voikko\footnote{https://voikko.puimula.org/}.

Omorfi \cite{pirinen2015development}\footnote{https://github.com/flammie/omorfi} is currently the most well supported tool for morphological analysis (including lemmatization) and generation. It is an FST (finite-state transducer) based tool developed on HFST (Helsinki finite-state technology) \cite{linden2013hfst} and it works together with constraint grammar (CG) based disambiguators and syntactic parsers available in the Giellatekno \cite{Moshagen2014} repositories\footnote{https://github.com/giellalt/lang-fin/tree/main/src/cg3}.

FinnPos\footnote{https://github.com/mpsilfve/FinnPos} \cite{silfverberg2016finnpos} is another morphological tagger and lemmatizer tool based on CRF (conditional random field). There have been recently more data driven approaches focusing on Finnish \cite{silfverberg-hulden-2018-initial}.

While rule-based tradition has been strong in the past\footnote{See \citealp{pirinen2019neural} for some comparison between rules and neural networks}, there are several machine learning driven dependency parsers for Finnish, such as the statistical one\footnote{https://turkunlp.org/Finnish-dep-parser/}  \cite{haverinen2013tdt} and neural one\footnote{http://turkunlp.org/Turku-neural-parser-pipeline/} \cite{udst:turkunlp} by TurkuNLP.

Out of the aforementioned tools Omorfi (and the CG disambigator) and the machine learning based parsers are available to use through a Python package named UralicNLP\footnote{https://github.com/mikahama/uralicNLP} \footnote{https://github.com/mikahama/uralicNLP/wiki/Dependency-parsing} \cite{uralicnlp_2019}.

As Finnish data is available in several multilingual datasets, there are many multilingual approaches for parsing \cite{qi-etal-2020-stanza}\footnote{https://stanfordnlp.github.io/stanza/} \cite{spacy}\footnote{https://spacy.io/} and morphology \cite{aharoni-goldberg-2017-morphological,nicolai-yarowsky-2019-learning,silfverberg-tyers-2019-data,gronroos2020morfessor}.

The fact that spoken Finnish is very different to standard Finnish has drawn some attention in the past \cite{jauhiainen-2001-using} and recently \cite{partanen-etal-2019-dialect}. The latter leading to a Python library called Murre\footnote{https://github.com/mikahama/murre} for automatic normalization of dialectal Finnish.

Non-standard data has been an issue in digital humanities (DH) projects \cite{makela2020wrangling}, and lately there have been efforts in automatically correcting OCR errors in existing historical datasets \cite{kettunen2015keep,drobac2020optical,1647b325b0f14e76b14c259d9a24c108,duong2020unsupervised}.

Named entity recognition has also been under study with FiNER\footnote{https://github.com/Traubert/FiNer-rules/blob/master/finer-readme.md} and its recently released data \cite{ruokolainen2019finnish}. There is also another recent BERT \cite{devlin-etal-2019-bert} based approach\footnote{https://turkunlp.org/fin-ner.html} to the topic \cite{luoma2020broad}.

There have been several approaches to language detection including detection of Finnish from web corpora (see  \citealp{jauhiainenjauhiainenjauhiainen}). Similarly, native Finnish has been automatically identified from learner's Finnish \cite{malmasi2014finnish}.

In summary, parsing has been researched on different levels of language such as syntax, morphology, POS and NER tagging, and lemmatization. It has been mainly focusing on standard well-formed Finnish, although there are methods for coping with dialectal Finnish and OCR errors as well. 

\subsection{Generation}

The lowest level of natural language generation is surface realization (see \citealp{reiter1994has}), and for that there are tools such as Omorfi and Syntax Maker\footnote{https://github.com/mikahama/syntaxmaker} \cite{hamalainen-rueter-2018-development}. The latter uses Omorfi for morphological inflection while it takes care of higher level morphosyntax such as case government and agreement.

There is a strong computational creativity focus in Helsinki and it also shows in Finnish NLG, as there are several poem generators such as Keinoleino\footnote{https://github.com/mikahama/keinoleino} \cite{hamalainen2018harnessing}, Poeticus \cite{toivanen2012corpus} and others \cite{hamalainen2019generating,hamalainen2019let}. There is also an interactive poem generator tool called \textit{Runokone} (Poem Machine)\footnote{http://runokone.cs.helsinki.fi/} \cite{hamalainen2018poem}.

Recently there have been several approaches to enhancing existing news headlines \cite{alnajjar2019no,ramo2021using}. And some approaches to generating entire news articles automatically \cite{kanerva2019template,haapanen2020recycling}.

Paraphrase generation \cite{sjoblom2020paraphrase} has also become a researched topic with the availability of monolingually aligned parallel corpora \cite{creutz2018open}. There is also an approach to converting standard Finnish text into different dialects \cite{hamalainen2020automatic}.

Finnish is a typical language for machine translation tasks and it is not uncommon to see it featured in several papers that deal with multiple languages. However, there are several papers that focus on Finnish in particular \cite{hurskainen2017rule,hamalainen2019template,pirinen-2019-apertium,tiedemann-etal-2020-fiskmo}.

There is also a recent approach to dialog generation in Finnish \cite{leino2020finchat}. Also non-native language learner's errors have been corrected successfully automatically \cite{creutz2019toward}.

To summarize the approaches, there are several generators for poetry and news that benefit from the available surface realizers. Paraphrasing, dialect adaptation, dialog generation and learners' error correction are domains with some research with potential for new discoveries in the future. Machine translation gets frequently attention from different researchers. There are several more NLG tasks (see \citealt{gatt2018survey}) that have not been researched at all in Finnish, which means that there is a lot of room for more research on this topic.

\subsection{Semantics}

Vector representations of meaning have become common place in NLP and Finnish is no exception with the availability of pretrained word2vec\footnote{http://vectors.nlpl.eu/repository/} \footnote{https://bionlp.utu.fi/finnish-internet-parsebank.html} \cite{laippala2014syntactic,kutuzov2017word} and fastText\footnote{https://fasttext.cc/docs/en/pretrained-vectors.html} \cite{bojanowski2017enriching} models.

BERT models have also become available as part of the multilingual BERT model\footnote{https://github.com/google-research/bert} \cite{devlin-etal-2019-bert} or trained separately for Finnish\footnote{http://vectors.nlpl.eu/repository/} \footnote{https://github.com/TurkuNLP/FinBERT} \cite{kutuzov2017word,virtanen2019multilingual}. Even Elmo models have been made available for Finnish\footnote{https://www.clarin.si/repository/xmlui/handle/11356/1277} \cite{ulvcar2020high}.

In addition to the standard vector-based representations of meaning, there is another statistical model called SemFi\footnote{https://github.com/mikahama/uralicNLP/wiki/Semantics-(SemFi,-SemUr)} \cite{hamalainen2018extracting}. The model is a relational database that captures semantic relations of words based on their syntactic co-occurencies.

Before the era of machine learning, there were two prominent projects for modeling meaning computationally which have been translated into Finnish WordNet \cite{linden2010finnwordnet} and FrameNet \cite{linden2019finntransframe}.

With the similar ideology to the hand crafted resources, there have been several different linked data projects in Finland representing semantics in structured ontologies \cite{hyvonen2006culturesampo,nyrkko2018building,thomas2018co,koho2019warsampo}. Many of the linked data projects are available on the Linked Data Finland website\footnote{https://www.ldf.fi/}.

There is a Python library called FinMeter\footnote{https://github.com/mikahama/finmeter} \cite{hamalainen2019let} that has some higher level semantic tools for Finnish such as metaphor interpretation, word concreteness analysis and sentiment analysis. Sentiment analysis for Finnish has also been studied later on\footnote{a dataset https://github.com/Helsinki-NLP/XED} \cite{ohman2020xed,8931724,linden2020finnsentiment}. There is also research on topic modeling methods \cite{ginter2009combining, hengchen2018comparing,loukasmaki2019eduskunnan}.

Finnish is well supported by traditional representations of semantics and latest vector based models. There is a vast amount of linked data resources in a variety of domains. Higher-level semantics such as metaphor interpretation and sentiment analysis also have received their share of research interest, although there are many more questions related to pragmatics and figurative language that have not been researched, such as sarcasm detection, multi-hop reasoning and fake news detection to name a few.

\subsection{Speech}

Apart from Finnish speech being supported by companies, there are some open-source tools that can synthesize Finnish. Festival\footnote{https://www.cstr.ed.ac.uk/projects/festival/} has a Finnish voice named Suopuhe\footnote{http://urn.fi/urn:nbn:fi:lb-20140730144}, and eSpeak-ng\footnote{https://github.com/espeak-ng/espeak-ng} can even generate IPA characters for Finnish.

There are several more modern approaches to speech recognition \cite{enarvi2017automatic,varjokallio2021morphologically} and speech synthesis \cite{2be1dd96f88b49c5be7f1a06a9f2d99a,raitio2014voice}. Although, speech synthesis has not gained much interest in the recent years.

There are several approaches to analyzing speech prosody \cite{virkkunen2018prosodic,simko2020}. There is also some work on detecting different accents in spoken Finnish \cite{behravan2013foreign,behravan2015factors} and named entity recognition \cite{porjazovski2020named}.

In summary, several approaches exist for speech processing in Finnish relating to recognition, accents and prosody. However, speech synthesis has received a surprisingly small amount of attention in the recent past. With the emergence of neural models, new research on synthesis could reach to potentially interesting new contributions.

\section{Discussion and Conclusions}

In this survey, we have gathered research conducted on different aspects of NLP. We have included links to models and code implementations for most of the research papers. It has been a pleasant thing to notice that not only Finnish NLP research exists but also it is often not conducted in a closed fashion, but the actual research outputs have been made openly available for a wider community of people even outside of academia. This is crucial for any language that is relatively small, like Finnish. If Finnish academics did not release their research, there would not be many other people in the world that would produce high-quality tools for Finnish.

Digital extinction is something that many endangered languages are facing right now (see \citealt{kornai2013digital}). Therefore, it is important to ensure that NLP resources become openly available for endangered Uralic languages as well. Availability itself is not enough, however, as the resources need to be easy to find and use. Despite the fact that we have open NLP tools for Finnish, we are still far a way from a world where machines use our language fluently. Finnair's in-flight entertainment system still announces happliy: \textit{*saavumme kohteeseen Helsinki} (*we arrive in destination Helsinki) instead of expressing it correctly, \textit{saavumme Helsinkiin} (we arrive in Helsinki), Google Doc's spell checker does not recognize mostly any inflectional form with a possessive suffix and predictive text in mobile keyboards suggest overly formal normative Finnish only.

While Finnish NLP has come far in terms of academic research and tools built as a result, we as a nation are still far away from having Finnish language technology fully integrated into the systems we use every day. Many of the problems have been solved already, it is just the matter of the industry finding out about the NLP tools that are out there.

We have limited our survey to NLP tools and methods only. We know that there are a plethora of language resources available for Finnish as well. Based on our experiences, many corpora are well hidden and digging them up is a time consuming effort worthy of a separate survey paper. Unfortunately the Finnish practice of describing data on Metashare\footnote{https://metashare.csc.fi/} is very unhelpful in this respect because the metadata descriptions in the service hardly ever contain information about where to access the data, how to cite it and who the real authors are.

% include your own bib file like this:
%\bibliographystyle{acl}
%\bibliography{acl2018}
\bibliography{acl2018}
\bibliographystyle{acl_natbib}

\end{document}